# Existence for Stable Rotating Star-Planet Systems

### v1

@@ 4 Existence of Constrained Minimizers of $E_{J}(\rho)$ on ${W}_{m}$
For $\lambda_{m,R}$, since the volume (measure) of $\Omega_{m}$ is ${Vol}:=\frac{4}{3}\pi\left(\frac{\eta}{4}\right)^{3}=$ $\frac{\pi J^{6}}{48m^{6}(1-m)^{6}}$, while the mass in $\Omega_{m}$ is $\int_{\Omega_{m}}\rho_{R}=m$. Claim: $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$. In fact, if not, then $\rho_{R}(x)-\frac{m}{Vol}>0$ almost everywhere in $\Omega_{m}$, while $\int_{\Omega_{m}}\left(\rho_{R}(x)-\frac{m}{Vol}\right)=0$ implies $\rho_{R}(x)-\frac{m}{Vol}=0$ almost everywhere in $\Omega_{m}$ [51, Section 4.3]. Since the measure of $\Omega_{m}$ is not 0, it makes a contradiction. Now we have $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$, which means there is at least a point $x_{0}$ in $\Omega_{m}$ such that $\rho_{R}\left(x_{0}\right)\leq\frac{m}{Vol}$ and (33) holds true. Notice ${Vol}\sim m^{-6}$ as $m\ll 1$, i.e. $\exists C_{1}>C_{2}>0$, such that if $m$ is sufficiently small, then $C_{2}m^{-6}\leq{Vol}\leq C_{1}m^{-6}$. Therefore, $\frac{m}{Vol}\sim$ $m^{7},A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\left(x_{0}\right)\stackrel{{\scriptstyle\gamma>\frac{3}{2}}}{{=}}O\left(m^{\frac{7}{3}}\right)$. By construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\forall y\in\Omega_{1-m},\left|x_{0}-y\right|\leq\operatorname{diam}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}},V_{\rho_{R}}\left(x_{0}\right)\geq$ $\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{\left|x_{0}-y\right|}d^{3}y\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)$ when $0<m<m_{1}$ for some $m_{1}$. Let $R>\frac{m_{1}}{{Vol}\left(m_{1}\right)}$, then $\rho_{R}(x_{0})$ that satisfies the above relation exists, and (33) holds true thus
@@
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall 0<m<M$ and $\forall R>0$, $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $I\left(\rho_{1-m}+\rho_{m}\right)>{Cm}^{-3}$.
@@
$\rho(m)=\rho_{1-m}+\rho_{m}\in{W}_{m,R}$. The centers of mass $\bar{x}\left(\rho_{1-m}\right)\in\Omega_{1-m},\bar{x}\left(\rho_{m}\right)\in\Omega_{m}$, as well as their projections onto $z=0$, are separated by at least $\frac{\eta}{2},\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$. Due to Lemma 2.4, we have
$I(\rho(m))=\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+I\left(\rho_{1-m}\right)+I\left(\rho_{m}\right)\geq\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)\geq\frac{J^{4}}{4\mu^{3}}$
Since $\mu=m(1-m)$, we know $\exists C>0$, such that $I(\rho(m))>Cm^{-3}$ when $m\rightarrow 0$. ∎
@@
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $0<m<M$, $R>0$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $v(x,m)=\frac{Jr(x-\bar{x}(\rho(m)))}{I(\rho(m))}<Cm$.
@@
We want to apply bootstrap method. Thanks to Corollary 4.10, when $m$ is sufficiently small, $\exists C_{1}>0$, such that $\forall x\in\Omega=\Omega_{m}\cup\Omega_{1-m},v^{2}(x)\leq C_{1}m^{2}$. Moreover, similarly as above if $x\in\Omega$, we have $\forall y\in\Omega_{1-m},|x-y|\leq\operatorname{diam}\left(\Omega_{m}\cup\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}},V_{\rho_{R}}(x)\geq$ $\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{|x-y|}d^{3}y\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>C_{2}m^{2}$ for some $C_{2}>0$ when $m$ is small enough. Therefore, we know $\exists\widetilde{C}>0$, such that $\forall x\in\Omega$, $v^{2}(x)\leq\widetilde{C}V_{\rho_{R}}(x)$. By (34), and Lemma 4.8 (Lagrange-multipliers are less then 0), when $R>R_{1}$, we have if $\rho_{R}(x)>0$, then $A^{\prime}\left(\rho_{R}(x)\right)\leq V_{\rho}(x)+v^{2}(x)\leq CV_{\rho}(x)$. If $\rho_{R}(x)=0$, the $0=$ $A^{\prime}\left(\rho_{R}(x)\right)\leq CV_{\rho_{R}}(x)$ is obvious.
Thanks to Lemma 4.7, we know if $R>R_{0}+R_{1},\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}}\leq k_{0}$. By HardyLittlewood-Sobolev Inequality (Proposition A.3), with an abuse of notation $C$, we have $\left\|V_{\rho_{R}}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}\leq$ $C\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)}\leq C$. $A^{\prime}\left(\rho_{R}\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\leq CV_{\rho}$ implies $\left\|\rho_{R}\right\|_{L^{12(\gamma-1)}}=$ $\left(\int\rho_{R}{}^{12(\gamma-1)}(x)d^{3}x\right)^{\frac{1}{12(\gamma-1)}}\leq C\left(\int V_{\rho_{R}}^{12}(x)d^{3}x\right)^{\frac{1}{12(\gamma-1)}}=C\left\|V_{\rho_{R}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C$. Since we know $\gamma>\frac{3}{2}>\frac{4}{3}$ and $\Omega$ is bounded, by interpolation inequality (Proposition A.2) we have $\left\|\rho_{R}\right\|_{L^{6}}\leq C$. Since we also know $\left\|\rho_{R}\right\|_{L^{1}\left(\mathbb{R}^{3}\right)}=1$ are bounded, thanks to Proposition A.4, we have $V_{\rho_{R}}$ is continuous and $\left\|V_{\rho_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is bounded uniformly. Again by $\rho_{R}{}^{\gamma-1}\leq CV_{\rho_{R}}$ we know $\left\|\rho_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq\hat{C}$ for some $\hat{C}>0$ when $R>R_{1}+R_{0}$. When $R\leq R_{1}+R_{0}$, by construction of $W_{m,R}$ we know $\left\|\rho_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq R_{1}+R_{0}$. Pick $k_{1}=\max\left\{\hat{C},R_{1}+R_{0}\right\}+1$ and then we get the result. For $R\geq k_{1}$, $E_{J}(\rho_{R})=E_{J}(\rho_{k_{1}})$ comes from the fact that they both minimize $E_{J}$ on $W_{m,k_{1}}$. ∎
@@
Given any $\sigma\in W_{m}\subset{L^{1}(\mathbb{R}^{3})\cap L^{\frac{4}{3}}(\mathbb{R}^{3})}$, the $L^{\infty}$ norm of $\sigma$ may be unbounded, hence in order to show the “double constrained minimizer” $\rho_{R}$ is indeed a constrained minimizer for $R$ large enough, we need to prove that energy of $\sigma$ still exceeds energy of $\rho_{R}$, which was not addressed in McCann’s work [46]. We provides a supplementary proof for it by adapting the arguments in the single-star case [4, Section 6].
@@
To prove the existence of constrained minimizer, let $k_{1}$ is given in Theorem 4.11, we have $E_{J}\left(\rho_{R}\right)=E_{J}\left(\rho_{k_{1}}\right)$ for $R\geq k_{1}$ due to Theorem 4.11. We will prove $\rho_{k_{1}}$ is a constrained minimizer. First we note that $\rho_{k_{1}}\in W_{m,k_{1}}\subset L^{1}(\mathbb{R}^{3})\cap L^{\infty}(\mathbb{R}^{3})$, then $E_{J}(\rho_{k})<\infty$ due to Remark 2.5. Given any $\sigma\in W_{m}\subset{L^{1}(\mathbb{R}^{3})\cap L^{\frac{4}{3}}(\mathbb{R}^{3})}$, due to Remark 2.2 we know $G(\sigma,\sigma)<\infty$. For $\sigma$ with $U(\sigma)=\int A(\sigma(x))d^{3}x=\infty$, we know $E_{J}(\rho_{k_{1}})<E_{J}(\sigma)=\infty$. For $\sigma$ with $U(\sigma)<\infty$, let
@@
When $R$ is large enough, we can find a $\tau\in\widetilde{P_{R}}(\sigma)$, with $\tau\geq 0$ and $\int_{\Omega_{m}}\tau=\int_{\Omega_{1-m}}\tau=1$. (See e.g. [15, Remark 4.10].) Let $\sigma_{s}^{[1]}(x),\sigma_{s}^{[2]}(x)$ and $\sigma_{s}$ be
@@
$\sigma_{s}^{[2]}(x)=\left\{\begin{array}[]{ll}\left(m-\int_{\Omega_{m}}\sigma_{s}^{[1]}\right)\tau,&\text{ if }x\in\Omega_{m}\\ \left((1-m)-\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\right)\tau,&\text{ if }x\in\Omega_{1-m}\\ 0,&\text{ otherwise }\end{array}\right$.
@@
By construction we have $\int_{\Omega_{m}}\sigma_{s}=m$ and $\int_{\Omega_{1-m}}\sigma_{s}=1-m$, then $\sigma_{s}\in W_{m}$. Similar again to [15, Remark 4.10], we have $\int_{\Omega_{m}}\sigma_{s}^{[1]}\rightarrow\int_{\Omega_{m}}\sigma=m$, $\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\rightarrow\int_{\Omega_{1-m}}\sigma=1-m$, therefore $\sigma_{s}^{[2]}\rightarrow 0$ almost everywhere in $\mathbb{R}^{3}$ as $s\rightarrow\infty$ due to $\tau\in L^{\infty}$. $\sigma\in W_{m}$ implies $\sigma$ is finite almost everywhere, thus $\sigma_{s}^{[1]}$ increases to $\sigma$ almost everywhere as $s\rightarrow\infty$. Then we know when $s\rightarrow\infty$, $\sigma_{s}\rightarrow\sigma$ a.e. in $\mathbb{R}^{3}$. Moreover, for almost every point $x\in\left\{x\in\mathbb{R}^{3}\mid\tau(x)=0\right\}$, $A\left(\sigma_{s}\right)=A\left(\sigma_{s}^{[1]}\right)$ increases to $A(\sigma)$ (except at points where $\tau=\infty$). For $\Omega_{>0}\coloneq\{x\in\mathbb{R}^{3}\mid\tau(x)>0\}$ (a subset of $\Omega$ thus bounded), $\sigma_{s}\rightarrow\sigma$ implies $A\left(\sigma_{s}\right)\rightarrow A(\sigma)$ a.e. as $s\rightarrow\infty$, and we also know $\left|\sigma_{s}\right|\leq\left|\sigma_{s}^{[1]}\right|+\left|\sigma_{s}^{[2]}\right|\leq\sigma+\|\tau\|_{L^{\infty}}$ a.e., which implies $A(\sigma_{s})\leq A(\sigma+\|\tau\|_{L^{\infty}})$ a.e. We claim $A(\sigma+\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$. In fact, since $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ is convex, we know

### v2

@@ 4 Existence of Constrained Minimizers of $E_{J}(\rho)$ on ${W}_{m}$
For $\lambda_{m,R}$, since the volume (measure) of $\Omega_{m}$ is ${Vol}:=\frac{4}{3}\pi\left(\frac{\eta}{4}\right)^{3}=$ $\frac{\pi J^{6}}{48m^{6}(1-m)^{6}}$, while the mass in $\Omega_{m}$ is $\int_{\Omega_{m}}\rho_{R}\,dx=m$.
Claim: $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$, where $\mu$ denotes the Lebesgue measure. In fact, if not, then $\rho_{R}(x)-\frac{m}{Vol}>0$ almost everywhere in $\Omega_{m}$, while $\int_{\Omega_{m}}\left(\rho_{R}(x)-\frac{m}{Vol}\right)\,dx=0$ implies $\rho_{R}(x)-\frac{m}{Vol}=0$ almost everywhere in $\Omega_{m}$ [28, Section 4.3]. Since the measure of $\Omega_{m}$ is not 0, it makes a contradiction. Now we have $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$, which means there is at least a point $x_{0}$ in $\Omega_{m}$ such that $\rho_{R}\left(x_{0}\right)\leq\frac{m}{Vol}$ and (34) holds true. Notice ${Vol}\sim m^{-6}$ as $m\ll 1$, i.e., $\exists C_{1}>C_{2}>0$, such that if $m$ is sufficiently small, then $C_{2}m^{-6}\leq{Vol}\leq C_{1}m^{-6}$. Therefore, $\frac{m}{Vol}\sim$ $m^{7},A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\left(x_{0}\right)\stackrel{{\scriptstyle\gamma>\frac{3}{2}}}{{=}}O\left(m^{\frac{7}{3}}\right)$. By construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\forall y\in\Omega_{1-m}$,
$\left|x_{0}-y\right|\leq\operatorname{diam}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}}$
$V_{\rho_{R}}\left(x_{0}\right)\geq\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{\left|x_{0}-y\right|}\,dy\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)$
when $0<m<m_{1}$ for some $m_{1}$. Let $R>\frac{m_{1}}{{Vol}\left(m_{1}\right)}$, then $\rho_{R}(x_{0})$ that satisfies the above relation exists, and (34) holds true thus
@@
Similarly, $\exists m_{2}>0,\forall 0<m<m_{2},\exists\widetilde{x_{0}}\in\Omega_{1-m}$, such that $\rho_{R}\left(\widetilde{x_{0}}\right)\leq\frac{1-m}{Vol}$, and since $\gamma>\frac{3}{2}$,
@@
Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall 0<m<M$ and $\forall R>0$, $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $I\left(\rho_{1-m}+\rho_{m}\right)>{Cm}^{-3}$.
@@
$\rho(m)=\rho_{1-m}+\rho_{m}\in{W}_{m,R}$. The centers of mass $\bar{x}\left(\rho_{1-m}\right)\in\Omega_{1-m}$, $\bar{x}\left(\rho_{m}\right)\in\Omega_{m}$, as well as their projections onto $z=0$, are separated by at least $\frac{\eta}{2}$, where $\eta=\frac{J^{2}}{\mu_{r}^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$. Due to Lemma 2.4, we have
$I(\rho(m))=\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+I\left(\rho_{1-m}\right)+I\left(\rho_{m}\right)\geq\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)\geq\frac{J^{4}}{4\mu_{r}^{3}}$
Since $\mu_{r}=m(1-m)$, we know $\exists C>0$, such that $I(\rho(m))>Cm^{-3}$ when $m\rightarrow 0$. ∎
@@
Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $0<m<M$, $R>0$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $v(x,m)=\frac{Jr(x-\bar{x}(\rho(m)))}{I(\rho(m))}<Cm$.
@@
We want to apply bootstrap method. Thanks to Corollary 4.10, when $m$ is sufficiently small, $\exists C_{1}>0$, such that $\forall x\in\Omega=\Omega_{m}\cup\Omega_{1-m}$, $v^{2}(x)\leq C_{1}m^{2}$. Moreover, similarly as above if $x\in\Omega$, we have $\forall y\in\Omega_{1-m}$, $|x-y|\leq\operatorname{diam}\left(\Omega_{m}\cup\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}}$, and
$V_{\rho_{R}}(x)\geq\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{|x-y|}\,dy\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>C_{2}m^{2}$
for some $C_{2}>0$ when $m$ is small enough. Therefore, we know $\exists\widetilde{C}>0$, such that $\forall x\in\Omega$, $v^{2}(x)\leq\widetilde{C}V_{\rho_{R}}(x)$. By (35), and Lemma 4.8 (Lagrange-multipliers are less than 0), when $R>R_{1}$, we have if $\rho_{R}(x)>0$, then $A^{\prime}\left(\rho_{R}(x)\right)\leq V_{\rho}(x)+v^{2}(x)\leq CV_{\rho}(x)$. If $\rho_{R}(x)=0$, then $0=$ $A^{\prime}\left(\rho_{R}(x)\right)\leq CV_{\rho_{R}}(x)$ is obvious.
Thanks to Lemma 4.7, we know if $R>R_{0}+R_{1},\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}}\leq k_{0}$. By HardyLittlewood-Sobolev Inequality (Proposition A.1), with an abuse of notation $C$, we have $\left\|V_{\rho_{R}}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}\leq$ $C\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)}\leq C$. And then $A^{\prime}\left(\rho_{R}\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\leq CV_{\rho}$ implies
$\left\|\rho_{R}\right\|_{L^{12(\gamma-1)}}=\left(\int_{\mathbb{R}^{3}}\rho_{R}{}^{12(\gamma-1)}(x)\,dx\right)^{\frac{1}{12(\gamma-1)}}\leq C\left(\int_{\mathbb{R}^{3}}V_{\rho_{R}}^{12}(x)\,dx\right)^{\frac{1}{12(\gamma-1)}}=C\left\|V_{\rho_{R}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C$
Since we know $\gamma>\frac{3}{2}>\frac{4}{3}$ and $\Omega$ is bounded, by interpolation inequality [6, Section 4.2] we have $\left\|\rho_{R}\right\|_{L^{6}}\leq C$. Since we also know $\left\|\rho_{R}\right\|_{L^{1}\left(\mathbb{R}^{3}\right)}=1$ are bounded, thanks to Proposition A.2, we have $V_{\rho_{R}}$ is continuous and $\left\|V_{\rho_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is bounded uniformly. Again by $\rho_{R}{}^{\gamma-1}\leq CV_{\rho_{R}}$ we know $\left\|\rho_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq\hat{C}$ for some $\hat{C}>0$ when $R>R_{1}+R_{0}$. When $R\leq R_{1}+R_{0}$, by construction of $W_{m,R}$ we know $\left\|\rho_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq R_{1}+R_{0}$. Pick $k_{1}=\max\left\{\hat{C},R_{1}+R_{0}\right\}+1$ and then we get the result. For $R\geq k_{1}$, $E_{J}(\rho_{R})=E_{J}(\rho_{k_{1}})$ comes from the fact that they both minimize $E_{J}$ on $W_{m,k_{1}}$. ∎
@@
Given any $\sigma\in W_{m}\subset{L^{1}(\mathbb{R}^{3})\cap L^{\frac{4}{3}}(\mathbb{R}^{3})}$, the $L^{\infty}$ norm of $\sigma$ may be unbounded, hence in order to show the “double constrained minimizer” $\rho_{R}$ is indeed a constrained minimizer for $R$ large enough, we need to prove that energy of $\sigma$ still exceeds energy of $\rho_{R}$, which was not addressed in McCann’s work [24]. We provide a supplementary proof for it by adapting the arguments in the single-star case [4, Section 6].
@@
To prove the existence of constrained minimizer, let $k_{1}$ be given in Theorem 4.11, we have $E_{J}\left(\rho_{R}\right)=E_{J}\left(\rho_{k_{1}}\right)$ for $R\geq k_{1}$ due to Theorem 4.11. We will prove $\rho_{k_{1}}$ is a constrained minimizer. First we note that $\rho_{k_{1}}\in W_{m,k_{1}}\subset L^{1}(\mathbb{R}^{3})\cap L^{\infty}(\mathbb{R}^{3})$, then $E_{J}(\rho_{k})<\infty$ due to Remark 2.5. Given any $\sigma\in W_{m}\subset{L^{1}(\mathbb{R}^{3})\cap L^{\frac{4}{3}}(\mathbb{R}^{3})}$, due to Remark 2.2 we know $G(\sigma,\sigma)<\infty$. For $\sigma$ with $U(\sigma)=\int_{\mathbb{R}^{3}}A(\sigma(x))\,dx=\infty$, we know $E_{J}(\rho_{k_{1}})<E_{J}(\sigma)=\infty$. For $\sigma$ with $U(\sigma)<\infty$, let
@@
When $R$ is large enough, we can find a $\tau\in\widetilde{P_{R}}(\sigma)$, with $\tau\geq 0$ and $\int_{\Omega_{m}}\tau\,dx=\int_{\Omega_{1-m}}\tau\,dx=1$. (See e.g. [10, Section 4].) Let $\sigma_{s}^{[1]}(x),\sigma_{s}^{[2]}(x)$ and $\sigma_{s}$ be
@@
$\sigma_{s}^{[2]}(x)=\left\{\begin{array}[]{ll}\left(m-\int_{\Omega_{m}}\sigma_{s}^{[1]}\,dx\right)\tau,&\text{ if }x\in\Omega_{m}\\ \left((1-m)-\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\,dx\right)\tau,&\text{ if }x\in\Omega_{1-m}\\ 0,&\text{ otherwise }\end{array}\right$.
@@
By construction we have $\int_{\Omega_{m}}\sigma_{s}\,dx=m$ and $\int_{\Omega_{1-m}}\sigma_{s}\,dx=1-m$, then $\sigma_{s}\in W_{m}$. Similar again to [10, Section 4], we have $\int_{\Omega_{m}}\sigma_{s}^{[1]}\,dx\rightarrow\int_{\Omega_{m}}\sigma\,dx=m$, $\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\,dx\rightarrow\int_{\Omega_{1-m}}\sigma\,dx=1-m$, therefore $\sigma_{s}^{[2]}\rightarrow 0$ almost everywhere in $\mathbb{R}^{3}$ as $s\rightarrow\infty$ due to $\tau\in L^{\infty}$. $\sigma\in W_{m}$ implies $\sigma$ is finite almost everywhere, thus $\sigma_{s}^{[1]}$ increases to $\sigma$ almost everywhere as $s\rightarrow\infty$. Then we know when $s\rightarrow\infty$, $\sigma_{s}\rightarrow\sigma$ a.e. in $\mathbb{R}^{3}$. Moreover, for almost every point $x\in\left\{x\in\mathbb{R}^{3}\mid\tau(x)=0\right\}$, $A\left(\sigma_{s}\right)=A\left(\sigma_{s}^{[1]}\right)$ increases to $A(\sigma)$ (except at points where $\tau=\infty$). For $\Omega_{>0}\coloneq\{x\in\mathbb{R}^{3}\mid\tau(x)>0\}$ (a subset of $\Omega$ thus bounded), $\sigma_{s}\rightarrow\sigma$ implies $A\left(\sigma_{s}\right)\rightarrow A(\sigma)$ a.e. as $s\rightarrow\infty$, and we also know $\left|\sigma_{s}\right|\leq\left|\sigma_{s}^{[1]}\right|+\left|\sigma_{s}^{[2]}\right|\leq\sigma+\|\tau\|_{L^{\infty}}$ a.e., which implies $A(\sigma_{s})\leq A(\sigma+\|\tau\|_{L^{\infty}})$ a.e. We claim $A(\sigma+\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$. In fact, since $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ is convex, we know